# Combining straight-line and map-based distances to investigate the connection between proximity to healthy foods and disease

### v1

@@ 4.7 Findings from the Piedmont Triad
As expected, the imputation analysis fell in many ways between the naive, gold standard, and complete case analyses. The imputation prevalence ratio estimates for both outcomes closely resembled those from the gold standard analysis. They were notably closer for diagnosed diabetes ($\widehat{PR}=0.998$, 95% CI: $0.984,1.013$) than the complete case analysis estimates. The four analysis methods were very similar for obesity ($\widehat{PR}=0.996$, 95% CI: $0.986,1.005$). As with the complete case, the wider confidence intervals from the imputation analysis both captured the null. Therefore, neither health outcome would be considered significantly associated with neighborhood proximity to healthy foods. Imputation’s efficiency gain over the complete case analysis was not seen in the Piedmont Triad data like in the simulations (Section 3). As discussed in Section 4.4, the errors in these data aligned with the moderate-to-severe simulation settings considered, which could explain some of the efficiency loss.
@@
The most notable difference for the non-spatial models was the smaller standard errors relative to the spatial models for either outcome and almost all analysis methods. Importantly, different conclusions would have been drawn from the non-spatial models, as the smaller standard errors led to narrower 95% CIs. In the gold standard analysis, both health outcomes could be considered significantly associated with proximity to healthy foods. Specifically, farther proximity (worse access) would be significantly associated with lower prevalences of diagnosed diabetes and obesity. These differences in conclusions demonstrate the importance of modeling spatial data accordingly, as ignoring the autocorrelation between neighboring census tracts could lead to incorrect conclusions about the relationships of interest. There could also be unmeasured confounding (e.g., due to socioeconomic differences), since these models did not adjust for additional covariates.
@@ 5 Conclusion
Eating healthy food is a critical determinant of healthy living, yet many communities have inadequate access to fresh, nutritious food. Given the public health ramifications of disparities in food access and their potential connection to disparities in health, accurately identifying low-access, high-risk communities is a priority. Quantifying access requires an extensive number of distance calculations (e.g., between neighborhood population centers and healthy food stores), and these calculations are either (i) computationally simple methods but inaccurate or (ii) accurate methods but computationally complex. This trade-off is solved by adopting a partial validation study design and then devising a multiple imputation for measurement error framework to overcome the missing data. In doing so, our approach offers improved accuracy in quantifying food access and modeling its impact on health with computational ease. The proposed methods make large-scale analyses of accurate distance-based access measures feasible by overcoming the computational hurdles. Therefore, the geographic scope of access studies can expand to answer questions and drive decision-making for larger communities.
@@
Imputation was preferable over other model-based methods here for a few reasons. First, the covariate imputation framework can be used with various outcome models, whereas maximum likelihood estimators (e.g., Tang et al., 2015) need to be re-derived for any model changes. This flexibility was demonstrated in the real data analysis, where spatial correlation between disease prevalences in bordering neighborhoods was detected and incorporated into the analysis model with ease. Second, under correct analysis and imputation model specification, imputation offers consistent estimates in Poisson regression. Regression calibration is only approximately unbiased for nonlinear outcome models, although it has been found to work fairly well Shaw et al. (2020).
@@
In simulations, imputation captures the benefits of the more accurate, map-based access data despite this data being largely missing. Notably, the simulations also demonstrate that the proposed methods offered unbiased, often highly efficient estimates under either additive or multiplicative error mechanisms, making the methods broadly applicable to other modeling approaches for access and health. With the naive analysis shown to be biased toward or away from the null, the importance of accurately measuring access was illustrated in simulations and the Piedmont Triad analysis.
@@
The Poisson regression model makes the assumption of equal dispersion (i.e., equal mean and variance) for the analysis model, which may not be reasonable in practice. Fortunately, it would be straightforward to adapt the multiple imputation procedure here to another model, like negative binomial regression, if over- or underdispersion were present. In the real data, imputation’s efficiency gains over the complete case analysis were not as clear as in the simulations, but its estimates remained closer to the gold standard. Another limitation related to the Piedmont Triad analysis is using disease counts $Y$ that are predictions from their own models. Therefore, though many of these socioeconomic variables may have been relevant to the analysis models here, additional covariates ${\mathbf{Z}}$ were not included since they were potentially already used to predict $Y$. Defining the origin for distance calculations from the census tracts posed one final challenge. A central point was used to measure access for all residents, smoothing over the gradient of access within a census tract.
@@
There are several interesting directions for future work. First, the subset of queried neighborhoods can be sampled based on any fully available information, providing an opportunity for strategic design. Perhaps existing optimal designs for Poisson regression (e.g., Wang et al., 2006) or measurement error in other models (e.g., Amorim et al., 2021; Lotspeich et al., 2024) could be adapted. Second, map-based driving distance was used to gauge neighborhood proximity. With map-based driving time as another metric of food access that is expensive to obtain, it could be interesting to impute time-based food access from straight-line measures instead. Third, neighborhood rates of disease could also be related to unhealthy foods access, or whichever type of food (healthy or unhealthy) is more accessible.
@@ Acknowledgements
The authors gratefully acknowledge the Andrew Sabin Family Center for Environment and Sustainability at Wake Forest University for a seed grant that supported this work.

### v2

@@ 4.7.4 Imputation analysis
In many ways, the imputation analysis fell between the naive, gold standard, and complete case analyses. The imputation estimates for both outcomes closely resembled those from the gold standard analysis and offered narrower confidence intervals than the complete case. For example, the estimated prevalence ratios for proximity to healthy foods in non-metropolitan neighborhoods were $\widehat{PR}_{X}=0.99$ for diabetes (95% CI: $0.98,1.01$) and $\widehat{PR}_{X}=1.00$ for obesity (95% CI: $0.99,1.01$). Still, we would not conclude that either health outcome was significantly associated with proximity to healthy foods among non-metropolitan neighborhoods.
The associations between proximity and the disease outcomes for metropolitan neighborhoods were similar to those for non-metropolitan ones, with $\widehat{PR}_{XZ}=1.00$ for diagnosed diabetes (95% CI: $0.97,1.03$) and $\widehat{PR}_{XZ}=1.00$ for obesity (95% CI: $0.98,1.01$). After adjusting for proximity, however, metropolitan neighborhoods were expected to have $17\%$ lower prevalence of diagnosed diabetes ($\widehat{PR}_{Z}=0.83$, 95% CI: $0.74,0.92$) and similar prevalence of obesity ($\widehat{PR}_{Z}=0.99$, 95% CI: $0.93,1.05$). Imputation’s efficiency gains were seen for all coefficients, not just those relying on $X$. Therefore, unlike the complete case, the imputation analysis led to the same conclusions about all coefficients and outcomes as the gold standard.
4.7.5 Spatial versus non-spatial models
The most notable difference for the non-spatial models was smaller standard errors relative to the spatial models for either outcome and almost all analysis methods. Larger standard errors were expected for the mixed-effects models, as they capture the spatial autocorrelation between neighboring census tracts through the added random intercepts. With smaller standard errors in the non-spatial models, the efficiency gains for imputation over the complete case analysis depended on the outcome and coefficient, which was not seen for the spatial models where the gains were universal.
@@ 4.7.5 Spatial versus non-spatial models
Importantly, different conclusions would have been drawn from the non-spatial models (using nearly all analysis methods), as the smaller standard errors assuming independence between census tracts led to narrower 95% CIs. For example, both health outcomes could be considered significantly associated with proximity to healthy foods based on the non-spatial gold standard analyses. Specifically, farther proximity (worse access) would be significantly associated with lower prevalences of diagnosed diabetes and obesity in non-metropolitan neighborhoods. For diagnosed diabetes, this same association would have been expected in metropolitan neighborhoods, while for obesity a stronger association would have been expected for metropolitan than non-metropolitan ones. These differences in conclusions demonstrate the importance of modeling spatial data accordingly, as ignoring the autocorrelation between neighboring census tracts could lead to incorrect conclusions about the relationships of interest.
@@ 5 Conclusion
Eating healthy food is a critical determinant of healthy living, yet many communities have inadequate access to fresh, nutritious food. Given the public health ramifications of disparities in food access and their potential connection to disparities in health, accurately identifying low-access, high-risk communities is a priority. Quantifying access requires an extensive number of distance calculations (e.g., between neighborhood population centers and healthy food stores), and these calculations are either (i) computationally simple methods but inaccurate or (ii) accurate methods but computationally complex. This trade-off is solved by adopting a partial validation study design and then devising a multiple imputation for measurement error approach to overcome the missing data. In doing so, our approach offers improved accuracy in quantifying food access and modeling its impact on health with computational ease. The proposed methods make large-scale analyses of accurate distance-based access measures feasible by overcoming the computational hurdles. Therefore, the geographic scope of access studies can expand to answer questions and drive decision-making for larger communities.
@@
Imputation was preferable over other model-based methods here for a few reasons. First, the covariate imputation framework can be used with various outcome models, whereas maximum likelihood estimators must be re-derived for any model changes. [65] This flexibility was demonstrated in the real data analysis, where spatial correlation between disease prevalences in bordering neighborhoods was detected and incorporated into the analysis model with ease. Second, under correct analysis and imputation model specification, imputation offers consistent estimates in Poisson regression. Regression calibration is only approximately unbiased for nonlinear outcome models, although it has been found to work fairly well. [66]
@@
In simulations and the application to the Piedmont Triad, imputation captured the benefits of the more accurate, map-based access measures despite these data being largely missing. Notably, the simulations demonstrated that the proposed methods offered unbiased, often highly efficient estimates under either additive or multiplicative error mechanisms, making the methods broadly applicable to other modeling approaches for access and health. In the real data application, the imputation estimates were close to the gold standard and accompanied by narrower confidence intervals than the complete case. With the naive analysis shown to be biased toward or away from the null, the importance of accurately measuring access was also illustrated.
@@
The Poisson regression model assumed equal dispersion (i.e., equal mean and variance) for the analysis model, which may not be reasonable in practice. Fortunately, it would be straightforward to adapt the multiple imputation procedure here to another model, like negative binomial regression, if over- or underdispersion were present. Another limitation stemmed from using the PLACES [57] disease counts $Y$ that are SAEs (i.e., predictions from their own models) in the Piedmont Triad analysis. Though many other socioeconomic variables could be relevant to the analysis models here (e.g., poverty), they were not included as covariates since they were potentially already used to predict $Y$. Defining the origin for distance calculations from the census tracts posed one final challenge. A central point (the population center) was used to measure access for all residents, smoothing over the gradient of access within a census tract.
@@
There are several interesting directions for future work. First, the subset of queried neighborhoods can be sampled based on any fully available information, providing an opportunity for strategic design. Perhaps existing optimal designs for Poisson regression [67] or measurement error in other models [68, 69] could be adapted. Second, map-based driving distance was used to gauge neighborhood proximity. With map-based driving time as another metric of food access that is expensive to obtain, it could be interesting to impute time-based food access from straight-line measures instead. Third, neighborhood rates of disease could also be related to unhealthy food access, or whichever type of food (healthy or unhealthy) is more accessible.
@@ Acknowledgements
The authors gratefully acknowledge the Andrew Sabin Family Center for Environment and Sustainability at Wake Forest University for a seed grant that supported this work. Computations were performed using the Wake Forest University (WFU) High Performance Computing Facility, a centrally managed computational resource available to WFU researchers including faculty, staff, students, and collaborators.